Give an example of a divergent sequence which has two convergent subsequences.


Justify your claim.

1
Expert's answer
2022-03-15T12:55:14-0400

Sequence "a_n = \sin(\pi n/2)" is divergent, since it oscillates between "-1, 0, 1" as "n\to\infty". However, if we take only even "n=2k", subsequence "a_{2k} = \sin(\pi k)" will converge to as "k\to \infty".

We can also take "n = 4k+1" and the subsequence "a_{4k+1} = \sin(2\pi k+\pi/2)" will converge to "1" as "k\to \infty".


Answer. "a_n = \sin(\pi n/2)".